Database last updated on Mon, 02 Feb 2026 03:40:49 +0000 by localhost

Instrument constant changes for the week before 2023307 (3 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 17 rows)

Measurements and tests summary.


_jday_loc_serial_skcaodedsedhgslumuvzs
307TORONTO#015Next schedule will be to2mk2um (now to2mk2um ZS= 98 ZS= 98),to2mk2um,Waiting until ZA = 98 now 94.3,Waiting until ZA = 98 now 96.1,Waiting until ZA = 98 now 97.8373725116787119
(1 sec)

Dead time (0 sec)

Nothing to report.

Standard lamp R6 (3 sec)

File nameCountSerialExpected (+/-20) Measured 30-day history
B30723.01516#01519901795